Simplify. Assume variables are nonnegative.
-3sqrt50
sqrt(18ab) * sqrt(6ab^3)
sqrt54/sqrt24
sqrt(1/7)
sqrt12 + 2sqrt27 - sqrt48
: Simplify. Assume variables are nonnegative.
-3sqrt50
sqrt(18ab) * sqrt(6ab^3)
sqrt54/sqrt24
sqrt(1/7)
sqrt12 + 2sqrt27 - sqrt48

Answer by jim_thompson5910(9217)

I'll do the first three to get you started.

1) -3*sqrt(50)=-3*5*sqrt(2)=-15*sqrt(2)


2) sqrt(18ab) * sqrt(6ab^3)=sqrt((18ab)(6ab^3))=sqrt(108a^2b^4)=6ab^2*sqrt(3)

3) sqrt(54)/sqrt(24)=(sqrt(54)/sqrt(24))(sqrt(24)/sqrt(24))=(sqrt(54)sqrt(24))/(sqrt(24)sqrt(24))=(sqrt(54)sqrt(24))/(24)=(3*sqrt(6)*2*sqrt(6))/(24)=(6*6)/24=36/24=3/2